OpenStudy (jotopia34):

How do you integrate: (3t^2+t+4)/t^3+t

OpenStudy (jotopia34):

\[\int\limits_{1}^{\sqrt{3}}\frac{ 3t^2+t+4 }{ t^3+t }dt\]

OpenStudy (tkhunny):

Correctly? What's Partial your Fraction plan?

OpenStudy (jotopia34):

I tried breaking it apart for integration by parts, tried u sub, none of it is working

OpenStudy (jotopia34):

Ill show you what I did

OpenStudy (abb0t):

factor out a t from the denomiator \[\frac{ A }{ t } + \frac{ Bx+C }{ t^2+1 } = 3t^2...\]

OpenStudy (tkhunny):

Of course, abb0t means \(Bt + C\). In any case, you'll probably have to cut this up after the partial fractions. \(\dfrac{Bt}{t^{2}+1} + \dfrac{C}{t^{2}+1}\). Now, THERE's your ArcTangent.

OpenStudy (tkhunny):

One degree less than the denominator. Since the denominator cannot be factored, it is what it is.

OpenStudy (tkhunny):

That's it. One degree less than the degree of the denominator, Excellent thinking it through!

OpenStudy (jotopia34):

I still do not get this. My class notes say, If x^2+bx+c is an irriducible quad. factor of q(x) and (x^2+bx+c)^n, the highest power dividiing of q(x). To this factor, assign the sum of n partial fractions. Repeat for each irriducible factor. This makes zero sense to me. Can you please help translate it for me. Then maybe I can figure out how to actually do it, because I am stuck

OpenStudy (tkhunny):

I can't really help you to sort through that.

OpenStudy (jotopia34):

To you tube I go....

hartnn (hartnn):

still need help with your class notes ? i may be able to explain it with an example...

OpenStudy (anonymous):

I can help drive this home.